OpenStudy (anonymous):

what is the lim --> 1 of : (1-x)tan[(pi*x)/2]

OpenStudy (turingtest):

(1-x)tan[(pi*x)/2] is defined at x=0

OpenStudy (anonymous):

sorry as x goes to 1

OpenStudy (turingtest):

ah, that makes it more interesting

OpenStudy (turingtest):

so first off, we want to split this into sin and cos notice that the sine limit can be taken immediately, then see what you have left

OpenStudy (turingtest):

still with me here?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

tell me if you can read this\[\lim_{x\to1}f(x)\]^^^^ ?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

great, can you try writing out what you get after changing tan to sin/cos ? we'll take it from there

OpenStudy (anonymous):

i did this on my notebook or should i write it here

OpenStudy (turingtest):

whatever is faster for you, attache the file or write it here

OpenStudy (turingtest):

I wouldn't have distributed the 1-x, but so far so good now notice that you can take the limits of the sin terms already, which leaves you with what?

OpenStudy (anonymous):

YES I GET 1 but on the cos i still get zero

OpenStudy (turingtest):

slow down, just do the sin terms first

OpenStudy (turingtest):

when the sin terms limits are taken what is left over? do NOT take the limit of the cosine term yet, and put it all over a common denominator

OpenStudy (anonymous):

i f i take only the limit of the sin i get 1-1=0

OpenStudy (turingtest):

no, if you take the limits of the sin terms on the paper you gave attached, you get\[\lim_{x\to1}\frac1{\cos{\pi x\over 2}}-\frac x{\cos{\pi x\over 2}}=\lim_{x\to1}{1-x\over\cos{\pi x\over 2}}\]

OpenStudy (anonymous):

i shouldn't have distributed to (1-x)

OpenStudy (turingtest):

I said that earlier, but we can always work it backwards, so no worries :) now two limits that are usually just memorized (because proof without l'hospital is geometric and somewhat painful)\[\lim_{x\to0}\frac{\sin x}x=1\]and\[\lim_{x\to1}{\cos x\over1- x}=0\]hopefully you already know these

OpenStudy (turingtest):

the game now is to rearrange what we have left so it looks like one of those two forms, then we can use those limits to get the anwer to our problem

OpenStudy (turingtest):

this is, admittedly, the tricky part. There may be a faster way to do it than what I did, but why don't you play with that for a bit and see if you can get anywhere. I'll help you more once you get stuck.

OpenStudy (anonymous):

wait a second please

OpenStudy (turingtest):

take your time

OpenStudy (anonymous):

i know (cosx-1)/x equals zero

OpenStudy (anonymous):

i am not sure about the what you have written cosx/(1-x)=0

OpenStudy (turingtest):

same thing as the second one I wrote

OpenStudy (turingtest):

either way, we are actually going to use the first one (well, at least I did)

OpenStudy (turingtest):

even though at first glance it looks more like the second, but there is a trig identity that gets us from sin to cos....

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

oh I'm sorry I did write that second one wrong :P\[\lim_{x\to0}{1-\cos x\over x}=0\] but as I said, we're not going to use it

OpenStudy (turingtest):

let me know if you want a hint

OpenStudy (anonymous):

should i multiply with cos(pix/2) up and down and than use the Half-Angle Identities for cos(x/2)

OpenStudy (turingtest):

not how I did it, but I can't say it won't work until I try it sounds harder than my way though

OpenStudy (anonymous):

ok tell me your way ,i will try this later

OpenStudy (turingtest):

let y=x-1 play with that for some time...

OpenStudy (turingtest):

sorry, let y=1-x

OpenStudy (anonymous):

i got lim of y..>0 y/[cos(pi(y+1)/2)]

OpenStudy (turingtest):

good, now distribute pi/2 in the cos argument

OpenStudy (turingtest):

like I said though, use y=1-x

OpenStudy (turingtest):

so should be a - sign there, my bad

OpenStudy (anonymous):

yeah cos((piy+y)/2)

OpenStudy (anonymous):

am i write about the part of lim y ...>0 ????

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

but try that substitution and distribution again with y=1-x

OpenStudy (anonymous):

lim y..>0 y/cos(pi/2 - ypi/2 )

OpenStudy (turingtest):

nice, now do you see a trig identity on the bottom yet?

OpenStudy (anonymous):

the Half-Angle Identity?

OpenStudy (turingtest):

simpler cos(x) becomes sin(x), and vice-versa, through a simple formula

OpenStudy (anonymous):

Cofunction Formulas ?

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

never seen that before .....so y/sin(ypi/2)

OpenStudy (anonymous):

we multiply with pi/2 two up and down and get 2/pi :)

OpenStudy (turingtest):

sorry I stepped away for a moment yes, right answer :) cofunction formulas are handy, you should look at a unit circle and convince yourself they are true. it's easy to see geometrically again, nice job, see ya!

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

i saw you are a moderator here

OpenStudy (turingtest):

hey hey

OpenStudy (anonymous):

this is my first time in this forum ,,it is very nice the way it works , getting live help but isn't this very time consuming , you are not charging anything besides the ads that flow around ???

OpenStudy (turingtest):

No, though you can sign up to be a special member of some type for 10$/month (which I have not done). I don't work for OpenStudy, I am just a long-time user who they elected to police the place. That is, they don't pay me hehe. OS is not a non-profit, but we also get a lot of grants from places like the Gates and Narional Science foundations

OpenStudy (turingtest):

they didn't even use to have ads when I started 3 years ago ;(

OpenStudy (anonymous):

i have seen a lot of website that charge a whole lot of money for help like this ,you do it for free.......even if they charge small amounts like 50 cents for simple problems like this people would be ready to pay small ammounts like that ,

OpenStudy (anonymous):

ok thank you for your help and the info

OpenStudy (turingtest):

Well, I've heard of such things, but those guys do the work for you mostly from what I've seen.The philosophy here is student teaching student. I'm glad you get a lot out of it, but beware, plenty of people will just spit out answers hoping for medals with no explanation. My job as a moderator is to curb that type of abuse. Feel free to tag me with anymore questions, see ya!

OpenStudy (anonymous):

great see you around

OpenStudy (anonymous):

i will asksometime and sometimes try to help someone else when i have time .....